OpenStudy (anonymous):

which is greater? (j=1 to n)summation(square(j)) or (j=1 to square(n))summation j

OpenStudy (unklerhaukus):

is this your question \[\sum\limits_{j=1}^nj^2\]or\[\sum\limits_{j=1}^{n^2}j\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you please answer it for me?

OpenStudy (unklerhaukus):

i think it depends on \(n\)

OpenStudy (anonymous):

i will go with door #2

OpenStudy (anonymous):

if i am not mistaken, first one is a polynomial of degree 3 and second one is a polynomial of degree 4, so it is larger then again i could be wrong

OpenStudy (anonymous):

If n=1 they are equal and if n>1 the second one is larger.

OpenStudy (anonymous):

ok.you're correct with the degrees, but even the coefficient matters....@satellite73

OpenStudy (anonymous):

you have actual formulas for each of these, so i t should be straightforward to compare

OpenStudy (anonymous):

how do you prove it?@sauravshakya i knw u cn do it by substituting values,but otherwise...?

OpenStudy (anonymous):

there are different degrees in each of them,so u cnt compare...@satellite73

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes, i claim a polynomial of degree 4 is larger than a polynomial of degree 3 for sure after some point, the fourth degree will be necessarily larger than the third degree

OpenStudy (anonymous):

if you want to find out at what point the fourth degree surpasses the third degree, i suppose you could be fancy, subtract the third degree from the fourth degree, take the derivative, and see over what interval the function is increasing

OpenStudy (anonymous):

\[4*2^{3}>2^{4}\]

OpenStudy (anonymous):

yes, that is true, so for some numbers \(4x^3>x^4\) but not in the long run

OpenStudy (anonymous):

ok so its only after a certain interval....so in the question either of them could be greater depending on the interval?

OpenStudy (anonymous):

if you subtract you get \[\frac{n^4}{2}-\frac{n^3}{3}-\frac{n}{6}\] and this is increasing on \((1,\infty)\)

OpenStudy (anonymous):

yes conceivably the cube could be greater than the fourth degree for some numbers, but not after a certain point. in your case that point is 1

OpenStudy (anonymous):

so door2 will be the correct answer?